How do I Reduce the following expression to its lowest terms?
4u^2-36/u^2-2u-15

Answer by asha(30)

4u^2-36/u^2-2u-15
Numerator= 4(u^2-9) =4(u-3)(u+3) difference of two squares
Denominator=u^2-5u+3u-15 splitting the middle term such that sum is -2
and product is 15.
=u(u-5)+3(u-5) taking common terms
=(u-5)(u+3)
the sum is 4(u+3)(u-3)/(u-5)(u+3)
cancelling common terms we get the answer
4(u-3)/(u-5)
good luck !!